OpenStudy (anonymous):

Choose the correct simplification of x6/x3.

OpenStudy (luigi0210):

\[\frac{ x^6 }{ x^3 }=x^(6-3)\]

OpenStudy (anonymous):

is 3

OpenStudy (luigi0210):

Dividing exponents is the same as subtracting them

OpenStudy (luigi0210):

Yes, it would be x^3

OpenStudy (anonymous):

thnx